How many cubed pieces of fudge that are 3 inches on an edge can be packed into a Christmas tin that is 9 inches deep by 12 inches wide by 8 inches high with the lid still being able to be closed? I got 32 but the answer is 24 and I don't know how they got it

Answer by Abbey(339)

The volume of the christmas tin is found by multiplying the depth by the width by the height:
8*8*12=864 cubic inches.
Each cube of fudge measures 3 by 3 by 3, so it's volume is found the same way:
3*3*3=27 cubic inches
Then just divide the tin's cubic inches by the fudge's cubic inches to see how many pieces can fit:
864/27=32 pieces